Use the last expression for y(x,t) to find an expression for the transverse velocity vy of a particle in the string on which the wave travels.

Use the last expression for y(x,t) to fi
Answer #1

© y(x,t) = Acos ( 2 ( x-VD)). Vy=v(x, t) = dylxst) yst (Acos (9* €2-) any A sin ( 1 (2-ve) The vale of sine is bound b/w 1&1